OpenStudy (anonymous):

how would you find the vertex of the parabola ? x^2-6x+8=y

Parth (parthkohli):

Complete the square. x^2 - 6x = y - 8 x^2 - 6x + 9 = y + 1 (x - 3)^2 = y + 1 (x - 3)^2 - 1 = y So, the vertex is (3,1)

Parth (parthkohli):

You add (b/2)^2 to both sides.

OpenStudy (anonymous):

first coordinate of vertex is always \(-\frac{b}{2a}\) in your case it is \(-\frac{-6}{2}=3\) second coordinate is what you get when you replace \(x\) by 3 in this case you get \(3^2-6\times 3+8=9-18+8=-9+8=-1\)

OpenStudy (anonymous):

and so vertex it \((3,-1)\)

Parth (parthkohli):

Oh no

Parth (parthkohli):

@satellite73 check out my smartscore hehe

OpenStudy (anonymous):

way to go!

Parth (parthkohli):

thanks!

Parth (parthkohli):

38 days

OpenStudy (anonymous):

amazing don't you ever sleep? must be a friend of saifoo

Parth (parthkohli):

Lol no I'm not saifoo's friend...and I didn't sleep last night

OpenStudy (anonymous):

btw vertex is still \((3,-1)\)

OpenStudy (anonymous):

your next job is to learn \(\LaTeX\) so you can write up answers nice nice

Parth (parthkohli):

I know Latex.

OpenStudy (anonymous):

oh ok

Parth (parthkohli):

\(\Huge \color{Green}{\text{Here we go, satellite73.}}\)

OpenStudy (anonymous):

lol

Parth (parthkohli):

;) your mom is from Palestine right?